Integration by Partial Fractions


∫dt/(1+sint+cost)


1
Expert's answer
2021-05-11T06:35:08-0400

∫dt/(1 + sint + cost ) ......................(1)




Let tan t/2 = z

Differentiating the above equation with respect to t, we have


(1/2 ) sec2 ( t/2 ) dt = dz


dt = 2 dz/( sec2 ( t/2 ) )


dt = 2 dz/( 1 + z2) .....................[∵ sec2 x = 1+ tan2 x]



Again by half angle formulae we have


sin t = 2 tan ( t/2) /( 1+ tan2 (t/2) )


cos t = ( 1- tan2 (t/2) ) / ( 1+ tan2 t/2) )


Substituting the values of sin t, cos t and dt from above in equation (1) we have


∫ [ 2 dz/ ( 1 + z2) ] / [ 1+ 2z/(1 + z2) +  [ ( 1- z2) / ( 1+ z2) ]

∫ [2 dz / ( 1 + z2) ] / [ (2 +2z) / ( 1 + z2) ]


∫ 2 dz /( 2+ 2z)


∫ dz /(1+z)


ln | 1+z| + c


Substituting the value of z from above we have


ln | 1+ tan (t/2) | + c